OpenStudy (anonymous):

solve for x 50/x=75/21 is it 25, 588, 108 or 112?

OpenStudy (anonymous):

it should be x=14

OpenStudy (anonymous):

its one of those its a multiple choice question

OpenStudy (anonymous):

50*1.5=75 x*1.5=21 -> x= 21/1.5= 14

OpenStudy (anonymous):

50/x = 75/21 -> 50/75 = x/21 -> 2*21/3 = x = 2*7 = 14

OpenStudy (anonymous):

thanks it is 14 sorry about the confusion